find the area of a 30-60-90 triangles with longest leg of 12

Answer by stanbon(75887)

find the area of a 30-60-90 triangles with longest leg of 12
----
The longest leg must be opposite the 60 degree angle.
-----
Let the shortest leg be "x".
Then the hypotenuse is "2x".
-----
Equation:
x^2 + 12^2 = (2x)^2
x^2 + 144 = 4x^2
3x^2 = 144
x^2 = 144/3
---
x = 12sqrt(3)/3
---
x = 4sqrt(3)
=================
Area = (1/2) leg*leg
Area = (1/2) 12*4sqrt(3)
--
Area = 24sqrt(3) square units
=================================
cheers,
Stan H.
====================